OpenStudy (anonymous):

Translate into mathematical symbols. Eight times three is less than or equal to ten times five.

OpenStudy (jamesj):

\[ 8 \times 3 \leq 10 \times 5 \]

OpenStudy (jamesj):

And it turns out that's a true mathematical statement, because indeed \[ 24 \leq 50 \]